OpenStudy (anonymous):

find the resistance of a piece if copper wire of length 20mm and diameter 0.5mm, take the resistivity to be 1.68 x 10 ^-8 ohms

OpenStudy (anonymous):

\[R =\rho \times( length \div Area)\]

OpenStudy (anonymous):

\[Area =( \pi \times \times diamater \times diameter)\div 4\]